OpenStudy (anonymous):

What is the solution to the system? Use the substitution method and show your work. y = 2x + 3, y = -1x + 6 (**Please walk me through this, I have no idea what to do**)

OpenStudy (anonymous):

Off the top of my head, you set the equations equal to one another, so \[2x+3=-1x+6\]

OpenStudy (anonymous):

(Please correct me if I'm wrong, the substitution method is unfamiliar to me XD)

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html This should help you. :) Step by Step.

OpenStudy (anonymous):

Thats right. @Green101

OpenStudy (amorfide):

you know what y is equal to, since both equations are equal to y, they are both equal to the same value therefore each equation is equal to the other which you can write as @Green101 stated 2x+3=-x+6 now solve for x, then substitute into one of the two given equations to work out y

OpenStudy (anonymous):

Mayne it won't work on urs. :/

OpenStudy (anonymous):

Maybe*

OpenStudy (anonymous):

Turn it into \[3x=3\] so \[x=1\]

OpenStudy (anonymous):

By collecting like terms

OpenStudy (anonymous):

Then plug that into one of the equations, \[y= (2)(1) + 3 \] so \[y=5\]

OpenStudy (anonymous):

okay i'm still not quite understanding it, i'm sorry :\ but thank you all for the help :)

pooja195 (pooja195):

ok lets make u understand :)

pooja195 (pooja195):

are you still here?

pooja195 (pooja195):

since they are both y equals you can plug either equation into anyone of there :)

pooja195 (pooja195):

*them

pooja195 (pooja195):

Ok this site is a bit laggy lol its taking a while for it to load so heres wht we can do y = 2x + 3 <----equation 1 y = -1x + 6 <------equation 2 Now you can choose which ever equation you want to plug in....to either equation :) im gonna plug in the 2nd equation into the first so i would sub in place of Y because the 2nd equation says Y equals y = -1x + 6 so this is how it would look: \[(-1x+6)=2x+3\]

pooja195 (pooja195):

from there all you need to do is solve subtract 6 from 3 which is -3 now your equation looks like this : \[-1x=2x-3\]

pooja195 (pooja195):

now subtract 2x from -1x and you get -3x your equation looks like this now : \[-3x=-3\] and now you divide by -3x since we want to isolate X

pooja195 (pooja195):

so you get 1 because -3/-3 is positive 1

pooja195 (pooja195):

x=1

pooja195 (pooja195):

then you take the one and plug it into one of the equations im gonna do the 2nd one \[y=-1(1)+6\] 1x1 = 1 +6 =7

pooja195 (pooja195):

y=7 Final answer : (1,7)

OpenStudy (anonymous):

@pooja195 okay sorry about that, i wasn't on bu I'm starting to understand it and I'm gonna try this a few times with the same equation and do it the way you did it up there and I'll just keep practicing. Thank you all so much for helping me :)

pooja195 (pooja195):

If you want we can pratice?